# Superconductivity in the repulsive Hubbard model on different geometries induced by density-assisted hopping

### v1

@@ Abstract
The theoretical description of high critical temperature superconductivity is a central subject in condensed matter physics. To tackle this problem, a commonly used model to try to describe the physics of cuprates is the two-dimensional Hubbard model Lee et al. (2006); Scalapino (2012); Arovas et al. (2022); Liu et al. (2025); Xiao et al. (2023); Qin et al. (2022). This is a very simple model with just two competing terms, yet it hosts a very rich phase diagram. Despite this very rich and complex phase diagram, it is still under investigation whether superconductivity occurs in the plain repulsive Hubbard model Qin et al. (2020, 2022). However, by adding an additional diagonal hopping $t^{\prime}$, superconductivity is enhanced Zheng et al. (2017); Qin et al. (2022); Xu et al. (2024); Zhang et al. (2025). Many extensions of the Hubbard model have been studied in the search for more realistic models of high-temperature superconductors Jiang and Devereaux (2019); Ponsioen et al. (2019); Jiang et al. (2020); Jiang and Kivelson (2021). These include density-assisted hopping terms Hirsch (1989); Hirsch and Marsiglio (1989); Simón and Aligia (1993); Japaridze and Müller-Hartmann (1994); Japaridze and Kampf (1999); Arrachea and Aligia (1999, 2000); Jiang et al. (2023); Controzzi (1999), which arise naturally from the down-folding of the three-band Hubbard model describing the Cu-O planes in cupratesVarma et al. (1987); Emery (1987) to a single-band Hubbard model. They were studied in the context of two-dimensional hole superconductivity to explain the asymmetry between particle and hole doping Hirsch (1989); Hirsch and Marsiglio (1989), and recently it has been shown numerically Jiang et al. (2023) that density-assisted hopping terms can increase the mobility of pairs of holes beyond mean-field effects. The inclusion of density-assisted hopping terms in the Hubbard model is motivated not only by solid-state systems, but also by possible direct experimental realizations in ultracold-atom experiments Jürgensen et al. (2014); Dutta et al. (2015); Meinert et al. (2016); Klemmer et al. (2024); Eckardt (2017); Görg et al. (2019); Chanda et al. (2025).
@@
We consider dimers formed by two sites with Hubbard repulsion. The two sites within the dimers are coupled by normal and density-assisted hopping terms. The inter-dimer coupling is a normal hopping. Examples of dimer lattices in one and two dimensions are the Hubbard ladder Noack et al. (1994, 1995, 1996, 1997); Dolfi et al. (2015); Shen et al. (2023); Sheikhan and Kollath (2020); Balents and Fisher (1996); Giamarchi (2004); Gannot et al. (2020) and the bilayer Hubbard model Lanatà et al. (2009); Maier and Scalapino (2011); Gall et al. (2021) (relevant for the description of superconductivity in nickelates Sun et al. (2023); Gupta et al. (2025)). In the ladder geometry [see Fig. 1(a)], the two chains are connected by normal and density-assisted hopping along the rungs, whereas in the bilayer, the two layers are connected this way.
@@
We show that density-assisted hopping can induce a superconducting phase (blue region in Fig. 1c) over an originally normal phase (yellow region in Fig. 1c). The origin of the superconductivity lies in an effective attractive interaction which is induced for a range of fillings by the density-assisted hopping (see Fig. 1b). Whereas this mapping to an effectively attractive model is independent of the underlying lattice geometry, we confirm our analytical results for the geometry of a ladder by the results from the numerically exact Matrix Product States (MPS) method (i.e., the density matrix renormalization group algorithm). We identify the effectively attractive interaction as the new pairing mechanism which is the origin of the induced superconducting correlations and thus the extended superconducting phase space.
@@
To obtain insights for large coupling of the dimers, it is useful to perform the transformation $\hat{c}_{k_{\perp},\sigma,j}=\left(\hat{c}_{1,\sigma,j}\pm\hat{c}_{2,\sigma,j}\right)/\sqrt{2},$ where the index $k_{\perp}=0,\pi$ corresponds to bands hereafter referred to as the bonding and antibonding bands, respectively. This transformation makes the perpendicular hopping term diagonal, $H_{\perp}=-t_{\perp}\sum_{\sigma,j}\left(\hat{n}_{0,\sigma,j}-\hat{n}_{\pi,\sigma,j}\right).$ For the non-interacting case ($U=0$), we find two independent bands shifted by an energy $2t_{\perp}$ where the bonding (antibonding) band is lower (higher) in energy (see sketch in Fig. 1b). For a filling $n<1$ and at $U=0$, both bands are partially occupied for $t_{\perp}<2t_{\parallel}|\cos(\pi n)|$, while for $t_{\perp}>2t_{\parallel}|\cos(\pi n)|$, only the bonding band remains partially filled due to increased band separation. For the remainder of this work, we will focus mainly on the regime where the two bands are well separated.
@@
Figure 2: Transition at $U=4t_{\parallel}$, $n=0.9375$ and $\tilde{t}_{\perp}=3t_{\parallel}$ for a 9696-rung ladder. (a) The gray and purple plots correspond to the central charge and the discontinuity of the momentum distribution $Z_{k_{F}}$, respectively. The orange plot corresponds to the inverse correlation length. (b) Momentum distribution for the up-spin in the bonding band, $n(k)$, for different values of the density-assisted hopping ratio ($c_{n}$). (c) Power-law exponents for the pair ($\eta_{d}$), density ($\eta_{n}$) and spin ($\eta_{sz}$) correlation functions. The plots are $1/\eta$, hence a larger value means than the correlations decay more slowly. The error bars were estimated by changing the fitting range and obtaining the maximum deviation. With translucid lines we plot the results for $L=32,48,64,80$ in panel (a), and $L=64,80$ in (c) (more opaque lines correspond to larger systems).
@@
We extract the central charge [see Fig. 2(a) (gray dots)] from the scaling of the entanglement entropy with subsystem size, i.e., $S_{\text{vN}}(l)=A+Bt(l)+c\ln d(l|L)/6$, where $d(l|L)=2L|\sin(\pi l/L)|/\pi$ is the conformal distance and $t(l)=\sum_{\alpha,\sigma}\mathrm{Re}\langle\hat{c}_{\alpha,\sigma,l}^{\dagger}\hat{c}_{\alpha,\sigma,l+1}^{\phantom{\dagger}}\rangle$ is the local kinetic energy (an example is given in Fig. 3 of the supplemental material). As expected, the central charge goes from $c\simeq 2$ in the C1S1 phase to $c\simeq 1$ in the C1S0 phase [Fig. 2(a)]. We attribute the deviations from the exact value to finite-size effects [different curve transparencies indicating the system size in Fig. 2(a)]. With increasing system size, one should recover a sharp step Läuchli and Kollath (2008).
@@
In order to test the sign change of the effective interaction in the bonding band obtained analytically, we numerically calculate the momentum distribution. It is obtained from the Fourier transform of the single-particle correlation function for the up-spin fermions, i.e., $n(k)=\text{FT}[\langle\hat{c}^{\dagger}_{0\uparrow i}\hat{c}^{\phantom{\dagger}}_{0\uparrow j}\rangle_{c}]$, where the subindex $c$ denotes the connected correlation function (see supplemental material for more details). We present the results for different density-assisted hopping ratio values in Fig. 2(b). For our calculations at $\tilde{t}_{\perp}=3t_{\parallel}$, the antibonding band is nearly empty, with approximately 97% of the particles occupying the bonding band. This observation is consistent with the picture of two well-separated bands, where only the bonding band is partially filled. We observe a jump at the Fermi momentum $k_{F}$ in the momentum distribution for all curves in Fig. 2(b) due to finite size effects. This allows us to define the momentum discontinuity as $Z_{k_{F}}=\max_{j}[n(k_{j})-n(k_{j+1})].$ Let us note that the jump for most parameters is due to finite-size effects, since in the thermodynamic limit and in the presence of interactions, $Z_{k_{F}}$ vanishes in one-dimensional systems due to the absence of well-defined quasiparticles. However, at the non-interacting point ($\tilde{U}=0$), i.e., at the transition point $c_{n}^{\star}$, for the effective one-band model (Eq. 2), the momentum distribution should be a step function and the value of $Z_{k_{F}}$ should reach unity. In Fig. 2(b), we show with orange lines the momentum distribution close to the critical point at $c_{n}=0.24$. It behaves as a sharp step function with $Z_{k_{F}}\simeq 0.97$. Furthermore, as shown in Fig. 2(a), $Z_{k_{F}}$ is not affected by system size at this point. We attribute the deviation from a step of $1$ to the remaining interactions with the antibonding band which has a very small occupation. The analytical picture of a sign change of the effective interaction is confirmed by our MPS results for $Z_{k_{F}}$ [see Fig. 2(a) (purple dots)]. The maximum of the momentum discontinuity $Z_{k_{F}}$ is estimated at $c_{n}^{*2}=0.23(1)$, agreeing with the value extracted from the gap opening $c_{n}^{*1}=0.24(1)$]. The precise scaling behavior of the momentum discontinuity $Z_{k_{F}}$ away from the critical point agrees with bosonization predictions, and the asymmetry with respect to this point is a consequence of different scaling exponents on either side (see supplemental material for more details).
@@
For low values of density-assisted hopping ratio $c_{n}$ [spin-gapless (C1S1) phase], we observe that the three exponents for spin, density, and pairs remain close to $\eta^{-1}\sim 0.5$. Among them, the pair correlations is the least dominant while the spin correlations appear to be slightly more dominant than the density correlations. These differences lie within our numerical resolution. Analytically, the long-distance decay of spin and density correlations is expected to be governed by the same Luttinger liquid exponent. We attribute the deviations from the expected behavior to finite-size effects.
@@
For intermediate-to-large values of the density-assisted hopping ratio the value of the inverse exponent associated with $d$-wave pair correlations increases rapidly, while the exponent corresponding to density correlations does not change significantly. The difference between the exponents increases well-beyond finite-size effects or the uncertainty coming from our fitting procedure. This shows the emergence of a very robust unconventional superconducting phase, where the pair correlations are considerably larger and decay more slowly than the other correlations in the system.
@@
At the transition point, i.e., the non-interacting point, all three exponents are expected to tend to the same value $\eta^{-1}=0.5$. We see this crossing in our data, up to small deviations from this predicted value of the order of the error bars (caused by finite system size and the limitations this imposes on our fitting procedure). The value of the density-assisted hopping ratio $c_{n}$ where the crossing takes place is $c_{n}^{\ast 3}=0.23(1)$. Again, this agrees well with the values obtained form the gap opening ($c_{n}^{\ast 1}$) and the maximum $Z_{k_{F}}$ discontinuity ($c_{n}^{\ast 2}$) up to the uncertainty coming from the discretization of our data points. It is also very remarkable that the numerically extracted value for the strong coupling is very close to the critical value predicted from Eq. 2 for weak interactions ($c_{n}^{\star}\simeq 0.26$).
@@
The same transition we discussed here for a specific value of the effective hopping ($\tilde{t}_{\perp}=3t_{\parallel}$) can be observed for a broader range of parameters. This is shown in Fig. 1(c) where we use the central charge in order to map out a phase diagram. The transition from the C1S1 Luttinger liquid phase to the C1S0 unconventional superconducting phase induced by the density-assisted hopping occurs over a wide parameter regime. The critical coupling predicted from Eq. 2 gives a good approximation for the observed transition (see purple dashed line in Fig. 1c).
@@
Summary and outlook.– We have studied the repulsive Hubbard model with density-assisted hopping in general geometries of dimerized lattices. We have shown that the density-assisted hopping can induce an effective attraction between holes that favors the formation of dimer-singlet pairs with a structure more complex than that of the $s$-wave pairs.
@@
We have presented strong numerical evidence that this picture, in principle valid for weak interactions, holds even in the presence of strong interactions in the ladder geometry. This supports the effective model, the attractive Hubbard model in the bonding band, which can also be used to predict the occurrence of superconducting phases with unconventional order in higher-dimensional geometries. In particular, in bilayer systems Lanatà et al. (2009); Maier and Scalapino (2011); Gall et al. (2021); Sun et al. (2023); Gupta et al. (2025) the density-assisted hopping between the planes would lead to an effective attractive two-dimensional Hubbard model for the bonding band, which is known to show superconductivity Scalettar et al. (1989); Gunnarsson et al. (2015); Qin et al. (2022). The $s$-wave superconductivity in the bonding band would lead to a non-local pairing between the bilayers.
@@
Acknowledgments: We thank J.B. Bernier and S. White for fruitful discussions. We acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 277625399-TRR 185 OSCAR (“Open System Control of Atomic and Photonic Matter”, B4), No. 277146847-CRC 1238 (“Control and dynamics of quantum materials”, C05), No. 511713970-CRC 1639 NuMeriQS (“Numerical Methods for Dynamics and Structure Formation in Quantum Systems”), and under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769. This research was supported in part by the National Science Foundation under Grants No. NSF PHY-1748958 and PHY-2309135. R.C. acknowledges support by the PNRR MUR project PE0000023 “National Quantum Science and Technology Institute” (NQSTI), through the cascade funding projects TOPQIN and SPUNTO.

### v2

@@ Abstract
The theoretical description of high critical temperature superconductivity is a central subject in condensed matter physics. To tackle this problem, a commonly used model to try to describe the physics of cuprates is the two-dimensional Hubbard model Lee et al. (2006); Scalapino (2012); Arovas et al. (2022); Liu et al. (2025); Xiao et al. (2023); Qin et al. (2022). This is a very simple model with just two competing terms, yet it hosts a very rich phase diagram. Despite this very rich and complex phase diagram, it is still under investigation whether superconductivity occurs in the plain repulsive Hubbard model Qin et al. (2020, 2022). However, the addition of diagonal hopping $t^{\prime}$, helps tip the balance in favor of superconductivity Zheng et al. (2017); Qin et al. (2022); Xu et al. (2024); Zhang et al. (2025). Many extensions of the Hubbard model have been studied in the search for more realistic models of high-temperature superconductors Jiang and Devereaux (2019); Ponsioen et al. (2019); Jiang et al. (2020); Jiang and Kivelson (2021). These include density-assisted hopping terms Hirsch (1989); Hirsch and Marsiglio (1989); Simón and Aligia (1993); Japaridze and Müller-Hartmann (1994); Japaridze and Kampf (1999); Arrachea and Aligia (1999, 2000); Jiang et al. (2023); Controzzi (1999), which arise naturally from the down-folding of the three-band Hubbard model describing the Cu-O planes in cupratesVarma et al. (1987); Emery (1987) to a single-band Hubbard model. They were studied in the context of two-dimensional hole superconductivity to explain the asymmetry between particle and hole doping Hirsch (1989); Hirsch and Marsiglio (1989). Recently it has been shown numerically Jiang et al. (2023) that the amplitude of the density-assisted hopping term in cuprates is 60% of that of the normal hopping and they can increase the mobility of pairs of holes beyond mean-field effects. The inclusion of density-assisted hopping terms in the Hubbard model is motivated not only by solid-state systems, but also by possible direct experimental realizations in ultracold-atom experiments Jürgensen et al. (2014); Dutta et al. (2015); Meinert et al. (2016); Klemmer et al. (2024); Eckardt (2017); Görg et al. (2019); Chanda et al. (2025).
@@
We consider dimers formed by two sites with Hubbard repulsion. The two sites within the dimers are coupled by normal and density-assisted hopping terms. The inter-dimer coupling is a normal hopping. Examples of dimer lattices in one and two dimensions are the Hubbard ladder Noack et al. (1994, 1995, 1996, 1997); Dolfi et al. (2015); Shen et al. (2023); Sheikhan and Kollath (2020); Balents and Fisher (1996); Giamarchi (2004); Gannot et al. (2020) and the bilayer Hubbard model Lanatà et al. (2009); Maier and Scalapino (2011); Gall et al. (2021). In the ladder geometry [see Fig. 1(a)], the two chains are connected by normal and density-assisted hopping along the rungs, whereas in the bilayer, the two layers are connected this way.
@@
We show that density-assisted hopping can induce a superconducting phase (blue region in Fig. 1c) over an originally normal phase (yellow region in Fig. 1c). The origin of the superconductivity lies in an effective attractive interaction which is induced for a range of fillings by the density-assisted hopping (see Fig. 1b). Whereas this mapping to an effectively attractive model is independent of the underlying lattice geometry, we confirm our analytical results for the geometry of a ladder by the results from the numerically exact Matrix Product States (MPS) method (i.e., the density matrix renormalization group algorithm). The amplitude of density-assisted hopping needed to induce superconductivity in our system is similar to the amplitude found for these terms in cuprates Jiang et al. (2023). We identify the effectively attractive interaction as the new pairing mechanism which is the origin of the induced superconducting correlations and thus the extended superconducting phase space.
@@
To obtain insights for large coupling of the dimers, it is useful to perform the transformation $\hat{c}_{k_{\perp},\sigma,j}=\left(\hat{c}_{1,\sigma,j}\pm\hat{c}_{2,\sigma,j}\right)/\sqrt{2},$ where the index $k_{\perp}=0,\pi$ corresponds to bands hereafter referred to as the bonding and antibonding bands, respectively. This transformation makes the perpendicular hopping term diagonal, $H_{\perp}=-t_{\perp}\sum_{\sigma,j}\left(\hat{n}_{0,\sigma,j}-\hat{n}_{\pi,\sigma,j}\right).$ For the non-interacting case ($U=0$), we find two independent bands shifted by an energy $2t_{\perp}$ where the bonding (antibonding) band is lower (higher) in energy (see sketch in Fig. 1b). For the remainder of this work, we will focus mainly on the regime where the two bands are well separated. For example, in the non-interacting ($U=0$) one-dimensional system below half filling ($n<1$), both bands are partially occupied for $t_{\perp}<2t_{\parallel}|\cos(\pi n)|$, while for $t_{\perp}>2t_{\parallel}|\cos(\pi n)|$, only the bonding band remains partially filled due to increased band separation.
@@
Figure 2: Transition at $U=4t_{\parallel}$, $n=0.9375$ and $\tilde{t}_{\perp}=3t_{\parallel}$ for a 9696-rung ladder. (a) The gray and purple plots correspond to the central charge and the discontinuity of the momentum distribution $Z_{k_{F}}$, respectively. The orange plot corresponds to the inverse correlation length. (b) Momentum distribution for the up-spin in the bonding and antibonding band, $n(k)$, for different values of the density-assisted hopping ratio ($c_{n}$). The occupation of the antibonding band stays below 3%. (c) Power-law exponents for the pair ($\eta_{d}$), density ($\eta_{n}$) and spin ($\eta_{sz}$) correlation functions. The plots are $1/\eta$, hence a larger value means than the correlations decay more slowly. The error bars were estimated by changing the fitting range and obtaining the maximum deviation. With translucid lines we plot the results for $L=32,48,64,80$ in panel (a), and $L=64,80$ in (c) (more opaque lines correspond to larger systems).
@@
We extract the central charge [see Fig. 2(a) (gray dots)] from the scaling of the entanglement entropy with subsystem size, i.e., $S_{\text{vN}}(l)=A+Bt(l)+c\ln d(l|L)/6$, where $d(l|L)=2L|\sin(\pi l/L)|/\pi$ is the conformal distance and $t(l)=\sum_{\alpha,\sigma}\mathrm{Re}\langle\hat{c}_{\alpha,\sigma,l}^{\dagger}\hat{c}_{\alpha,\sigma,l+1}^{\phantom{\dagger}}\rangle$ is the local kinetic energy (an example is given in Fig. 3 of the supplemental material). As expected, the central charge decreases from $c\simeq 2$ in the C1S1 phase to $c\simeq 1$ in the C1S0 phase [Fig. 2(a)]. We attribute the deviations from the exact value to finite-size effects [different curve transparencies indicating the system size in Fig. 2(a)]. With increasing system size, one should recover a sharp step Läuchli and Kollath (2008).
@@
In order to test the sign change of the effective interaction in the bonding band obtained analytically, we numerically calculate the momentum distribution. It is obtained from the Fourier transform of the single-particle correlation function for the up-spin fermions, i.e., $n_{k\perp}(k)=\text{FT}[\langle\hat{c}^{\dagger}_{k\perp,\uparrow,i}\hat{c}^{\phantom{\dagger}}_{k\perp,\uparrow,j}\rangle_{c}]$, where the subindex $c$ denotes the connected correlation function (see supplemental material for more details) and $k_{\perp}=0,\pi$ refers to the bonding and antibonding band, respectively. We present the results for different density-assisted hopping ratio values in Fig. 2(b). For our calculations at $\tilde{t}_{\perp}=3t_{\parallel}$, the antibonding band (small circles) is nearly empty with less than 3% of all fermions. However, the bonding band (big circles) represents approximately 97% of the particles. This observation is consistent with the picture of two well-separated bands, where only the bonding band is partially filled. We observe a jump at the Fermi momentum $k_{F}$ in the momentum distribution for all curves in Fig. 2(b) due to finite size effects. This allows us to define the momentum discontinuity as $Z_{k_{F}}=\max_{j}[n(k_{j})-n(k_{j+1})].$ Let us note that the jump for most parameters is due to finite-size effects, since in the thermodynamic limit and in the presence of interactions, $Z_{k_{F}}$ vanishes in one-dimensional systems due to the absence of well-defined quasiparticles. However, at the non-interacting point ($\tilde{U}=0$), i.e., at the transition point $c_{n}^{\star}$, for the effective one-band model (Eq. 2), the momentum distribution should be a step function and the value of $Z_{k_{F}}$ should reach unity. In Fig. 2(b), we show with orange lines the momentum distribution close to the critical point at $c_{n}=0.24$. It behaves as a sharp step function with $Z_{k_{F}}\simeq 0.97$. Furthermore, as shown in Fig. 2(a), $Z_{k_{F}}$ is not affected by system size at this point. We attribute the deviation from a step of $1$ to the remaining interactions with the antibonding band which has a very small occupation. The analytical picture of a sign change of the effective interaction is confirmed by our MPS results for $Z_{k_{F}}$ [see Fig. 2(a) (purple dots)]. The maximum of the momentum discontinuity $Z_{k_{F}}$ is estimated at $c_{n}^{*2}=0.23(1)$, agreeing with the value extracted from the gap opening $c_{n}^{*1}=0.24(1)$]. The precise scaling behavior of the momentum discontinuity $Z_{k_{F}}$ away from the critical point agrees with bosonization predictions, and the asymmetry with respect to this point is a consequence of different scaling exponents on either side (see supplemental material for more details).
@@
For low values of density-assisted hopping ratio $c_{n}$ [spin-gapless (C1S1) phase], we observe that the three exponents for spin, density, and pairs remain close to $\eta^{-1}\sim 0.5$. Among them, the pair correlations is the least dominant while the spin correlations appear to be slightly more dominant than the density correlations. These differences lie within our numerical resolution. Analytically, the long-distance decay of spin and density correlations is expected to be governed by the same Luttinger liquid exponent. Additionally, the bosonization predicts the relation $\eta_{n}\eta_{d}=1$ for long distances. We have studied the deviations from the expected behavior in the supplemental material and show that we can attribute them to finite-size effects.
@@
For intermediate-to-large values of the density-assisted hopping ratio, the value of the inverse exponent associated with $d$-wave pair correlations increases rapidly, while the exponent corresponding to density correlations does not change significantly. The difference between the exponents increases well-beyond finite-size effects or the uncertainty coming from our fitting procedure. This shows the emergence of a very robust unconventional superconducting phase, where the pair correlations are considerably larger and decay more slowly than the other correlations in the system.
@@
At the transition point, i.e., the non-interacting point, all three exponents are expected to tend to the same value $\eta^{-1}=0.5$. We see this crossing in our data, up to small deviations from this predicted value of the order of the error bars (caused by finite system size and the limitations this imposes on our fitting procedure). The value of the density-assisted hopping ratio $c_{n}$ where the crossing takes place is $c_{n}^{\ast 3}=0.23(1)$. Again, this agrees well with the values obtained from the gap opening ($c_{n}^{\ast 1}$) and the maximum discontinuity $Z_{k_{F}}$ ($c_{n}^{\ast 2}$) up to the uncertainty arising from the discretization of our data points. It is also very remarkable that the numerically extracted value for the strong coupling is very close to the critical value predicted from Eq. 2 for weak interactions ($c_{n}^{\star}\simeq 0.26$).
@@
The same transition we discussed here for a specific value of the effective hopping ($\tilde{t}_{\perp}=3t_{\parallel}$) can be observed for a broader range of parameters. This is shown in Fig. 1(c) where we use the central charge to map out the phase diagram. The transition from the C1S1 Luttinger liquid phase to the C1S0 unconventional superconducting phase induced by the density-assisted hopping occurs over a wide parameter regime. The critical coupling predicted from Eq. 2 gives a good approximation for the observed transition (see purple dashed line in Fig. 1c).
Thus, we have seen that in the dimerized ladder system, the density-assisted hopping induces superconductivity. Remarkably, the values at which the system becomes superconducting are similar to the values of density-assisted hopping estimated for cuprates Jiang et al. (2023). This transition is very robust, and we find its signatures in a much broader parameter regime than the stringent validity regime of the effective model (see also the supplemental material for higher interaction strength and other fillings).
The results we obtained are not restricted to the case of the ladder geometry. In the cases of dimerized structures in two and three dimensions, the corresponding effective models are the two- and three-dimensional Hubbard models, respectively. These models are known to host a superconducting phase for attractive interactions Qin et al. (2022); Scalettar et al. (1989); Gunnarsson et al. (2015); Keller et al. (2001); Toschi et al. (2005); Bauer et al. (2009). Thus, also in higher dimensions, the density-assisted hopping term induces a superconducting phase at weak coupling. As in the one-dimensional setup, we expect that the superconducting phase exists to larger interaction strength much beyond the assumed validity regime of the effective model.
@@
Summary and outlook.– We have studied the repulsive Hubbard model with density-assisted hopping on general geometries of dimerized lattices. We show that density-assisted hopping can induce interactions with a band-selective sign change. In the lower band, it generates an effective attraction between holes that favors the formation of dimer-singlet pairs with a structure more complex than conventional $s$-wave pairing. The resulting effective attraction therefore arises from a mechanism distinct from previously identified pairing mechanisms in Hubbard systems, and provides a route to emergent multiband superconductivity driven by correlated hopping. We have presented strong numerical evidence that this picture, in principle valid for weak interactions, holds even in the presence of strong interactions in the ladder geometry.
@@
This supports the effective model, the attractive Hubbard model in the bonding band, which can also be used to predict the occurrence of superconducting phases with unconventional order in higher-dimensional geometries. In particular, in bilayer systems Lanatà et al. (2009); Maier and Scalapino (2011); Gall et al. (2021); Sun et al. (2023); Gupta et al. (2025) the density-assisted hopping between the planes would lead to an effective attractive two-dimensional Hubbard model for the bonding band, which is known to show superconductivity Scalettar et al. (1989); Gunnarsson et al. (2015); Qin et al. (2022). The $s$-wave superconductivity in the bonding band would lead to a non-local pairing between the bilayers. The two-band bilayer Hubbard model is relevant for the description of the nickelate La3Ni2O7 Sun et al. (2023); Gupta et al. (2025); Maier et al. (2026), where the interlayer hopping is the largest and the NiO planes form a Lieb lattice, analogous to the CuO planes in cuprates. It would be worthwhile to consider if downfolding that model to a single band model can give rise to a significant density assisted hopping.
@@
Acknowledgments: We thank J.B. Bernier, T. Giamarchi and S. White for fruitful discussions. We acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 277625399-TRR 185 OSCAR (“Open System Control of Atomic and Photonic Matter”, B4), No. 277146847-CRC 1238 (“Control and dynamics of quantum materials”, C05), No. 511713970-CRC 1639 NuMeriQS (“Numerical Methods for Dynamics and Structure Formation in Quantum Systems”), and under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534769. This research was supported in part by the National Science Foundation under Grants No. NSF PHY-1748958 and PHY-2309135. R.C. acknowledges support by the PNRR MUR project PE0000023 “National Quantum Science and Technology Institute” (NQSTI), through the cascade funding projects TOPQIN and SPUNTO.